OpenStudy (anonymous):

how would you solve the following math problem 7 7/15-4 4/15 step by step answer is 16/5

OpenStudy (heisenberg):

You'll have to change this to full fractions (no whole numbers). Here is how you turn what you have (whole numbers + fractions) into just a fraction:

OpenStudy (heisenberg):

$$7 \frac{7}{15} = \frac{7 * 15 + 7}{15}$$

OpenStudy (heisenberg):

so doing that for each, it will be easy do simple subtraction...

OpenStudy (anonymous):

i thank you so much so after that i just have the answer of 16/5?

OpenStudy (heisenberg):

$$\frac{7*15+7}{15} - \frac{15*4 + 4}{15} = \frac{112}{15} - \frac{64}{15} = \frac{48}{15} = \frac{16}{5}$$

OpenStudy (heisenberg):

$$\frac{48}{15} = \frac{16}{5} (reduced)$$

OpenStudy (anonymous):

thank you so much for your help

OpenStudy (shadowfiend):

Btw, you reduce it by seeing that 5 is the denominator you want, and you have to divide 15 by 3 to get 5, so dividing both 48 and 15 by 3 gets you 16 / 5.

OpenStudy (anonymous):

thank you

OpenStudy (loki):

great work here!

OpenStudy (anonymous):

plz solve this 4 me g is inverse of f f(4)=5,f\[\prime\](4)=2/3,find g\[\prime\](5)

OpenStudy (anonymous):

@toni32: if using big numbers is too complicated, you can still use this method. 1st step: subtract the whole numbers: 7 - 4 = 3 2nd step: subtract the fraction: 7/15 - 4/15 directly as they have the same denominator.Ans: 3/15 reducing it to 1/5 3rd step: combine( or add) the whole number with the fraction: 3 + 1/5 = 3 1/5 (as a mixed fraction) or 16/5 (as an improper fraction). From arwin